solve for x log3 (2x-1)=2+log3 (x-2)
Answer by lwsshak3(11628)

solve for x log3 (2x-1)=2+log3 (x-2)
**
log3 (2x-1)=2+log3 (x-2)
log3 (2x-1)-log3 (x-2)=2
place under single log using division rule for logs
log3[(2x-1)/(x-2)]=2
convert to exponential form: base(3) raised to log of number(2)=number(2x-1)/(x-2)
3^2=(2x-1)/(x-2)=9
2x-1=9x-18
7x=17
x=17/7